# Beyond RMSE and MAE: Introducing EAUC to unmask hidden bias and unfairness in dyadic regression models

### v4

@@ II-B State-of-the-art evaluation: RMSE and Netflix Prize
III Motivation: Limitations of RMSE in Dyadic regression tasks
In this section, we demonstrate the limitations of using global error metrics like RMSE or MAE as the sole means for evaluating performance in dyadic regression tasks. As traditionally used, they evaluate considering the dataset as a whole without distinguishing between different users, items, user-item pairs or their associated traits. Therefore, while suitable for classic regression tasks, they may not sufficiently capture the intricacies of dyadic data.
To exemplify the limitations of RMSE we turn to a practical example using the Netflix Prize data [11] and a Matrix Factorization (MF) model [18], a state-of-the-art dataset and model also described later in Section V.
@@ III Motivation: Limitations of RMSE in Dyadic regression tasks
IV Bias-aware Evaluation of Dyadic Regression
This section introduces EAUC, our novel alternative metric that can detect and quantify eccentricity bias and its subsequent unfairness in dyadic regression models’ predictions by providing a holistic view of their performance: a good model should excel not only when predicting near the $DMV_{ui}$ of the dyad (the expected value for the (u,i) interaction) but also in cases with higher eccentricity, where the model really proves it has understood $u$’s and $i$’s latent characteristics rather than having memorized $\bar{r}_{u}$ and $\bar{r}_{i}$. Such eccentric cases are equally essential, often representing critical edge cases where fairness implications are highest.
IV-A Eccentricity-Area Under the Curve (EAUC)
The underlying principle of EAUC is to, by accounting for the eccentricity of examples, ensure that the measured performance of a model can reflect its ability to handle cases that deviate significantly from the average past values of the entities involved. This overcomes the main limitation of global error metrics like RMSE or MAE, where the performance in these uncommon cases has a negligible effect on the measured error.
EAUC can be computed through the following process, also codified in Algorithm 1:
@@ IV-A Eccentricity-Area Under the Curve (EAUC)
$EAUC(f,\mathcal{D})=\frac{\sum\limits_{x_{i}\in\mathcal{D}_{test}}\frac{(Ecc_{x_{i}}-Ecc_{x_{i-1}})*(\epsilon_{x_{i}}+\epsilon_{x_{i-1}})}{2}}{(max(r_{ui})-min(r_{ui}))^{2}}\quad$ (5) $\text{Where}\quad\quad\quad Ecc_{\{u,i,r_{ui}\}}=$ $|r_{ui}-\frac{1}{2}(\bar{r}_{u}+\bar{r}_{i})|,$ $\epsilon_{\{u,i,r_{ui}\}}=$ $(r_{ui}-f(u,i))^{2},$ $\bar{r}_{u}=$ $\frac{1}{|\mathcal{I}_{u}|}\sum_{i\in\mathcal{I}_{u}\subseteq\mathcal{I}_{\mathcal{D}_{train}}}r_{ui},$ $\bar{r}_{i}=$ $\frac{1}{|\mathcal{U}_{i}|}\sum_{u\in\mathcal{U}_{i}\subseteq\mathcal{U}_{\mathcal{D}_{train}}}r_{ui},$ $x_{i}\in$$\{u,i,r_{ui}\};$ $\text{With}\quad(Ecc_{x_{i}}-Ecc_{x_{i-1}})\geq$ 00 0
IV-B Interpretation of EAUC for model evaluation
@@
Moreover, since eccentricity bias leads to discriminatory outcomes when fairness is involved, EAUC also behaves as measure of fairness in dyadic regression tasks, being able to complement and synergize with classic fairness metrics. In this aspect, it is plausible that EAUC will become a reliable detector of an -until now- unknown source of unfairness in this type of task, i.e. eccentricity bias. As an indicator of unfairness in dyadic regression, EAUC may be used to directly detect the presence of eccentricity-based individual or group unfairness [19], which can later be characterized using utility or parity-based fairness metrics [20]. Conversely, the inverse is also possible: while most group or individual unfairness are often attributed to representational or class-stereotypical biases in the data [21], they often cannot fully explain the source of fairness[22]; EAUC and its analysis may be used to better characterize the source of otherwise opaque unfairness phenomenons. One final property of EUAC is its ability to, unlike classic individual and group-based metrics modelling parity and utility [23], detect eccentricity-based unfairness that affects all entities (e.g. users) of the population as a whole, rather than only specific subgroups or individuals. As a result, EAUC can improve our ability to both detect and better characterize new and known unfairness in dyadic regression.

### v5

@@ II-B State-of-the-art evaluation: RMSE and Netflix Prize
III The need for EAUC: Limitations of RMSE in Dyadic regression tasks
In this section, we justify the proposal of our bias-aware evaluation metric EAUC by experimentally proving the limitations of global error metrics like RMSE in dyadic regression tasks. As traditionally used, these metrics evaluate over the dataset as a whole without distinguishing between different users, items, user-item pairs or their associated traits, and therefore we novelly postulate that they do not sufficiently capture the intricacies of dyadic data. In the following, we reason and demonstrate the limitations of RMSE using the Netflix Prize data [11] and a Matrix Factorization (MF) model [18], a state-of-the-art dataset and model in the task, respectively.
@@ III The need for EAUC: Limitations of RMSE in Dyadic regression tasks
IV EAUC: Bias-aware Evaluation of Dyadic Regression
This section formalizes EAUC (Eccentricity-Area Under The Curve), our proposed novel alternative metric that can detect and quantify eccentricity bias and its subsequent unfairness in dyadic regression models’ prediction. In the following, we discuss both its theoretical foundation and its practical usage in evaluation policies.
IV-A Theoretical derivation of EAUC
The underlying theory of EAUC is to, by accounting for the eccentricity of examples, ensure that the measured performance of a model can reflect its ability to handle cases that deviate significantly from the average past values $DMV_{ui}$ of the entities involved; these cases often represent critical edge cases where fairness implications are highest, but they are neglected by global error metrics like RMSE. As codified in Algorithm 1, EAUC is computed as follows:
@@ IV-A Theoretical derivation of EAUC
$EAUC(f,\mathcal{D})=\frac{\sum\limits_{x_{i}\in\mathcal{D}_{test}}\frac{(Ecc_{x_{i}}-Ecc_{x_{i-1}})*(\epsilon_{x_{i}}+\epsilon_{x_{i-1}})}{2}}{(max(r_{ui})-min(r_{ui}))^{2}}\quad$ (5)
$\text{Where}\quad\quad\quad Ecc_{\{u,i,r_{ui}\}}=$ $|r_{ui}-\frac{1}{2}(\bar{r}_{u}+\bar{r}_{i})|,$
$\epsilon_{\{u,i,r_{ui}\}}=$ $(r_{ui}-f(u,i))^{2},$
$\bar{r}_{u}=$ $\frac{1}{|\mathcal{I}_{u}|}\sum_{i\in\mathcal{I}_{u}\subseteq\mathcal{I}_{\mathcal{D}_{train}}}r_{ui},$
$\bar{r}_{i}=$ $\frac{1}{|\mathcal{U}_{i}|}\sum_{u\in\mathcal{U}_{i}\subseteq\mathcal{U}_{\mathcal{D}_{train}}}r_{ui},$
$x_{i}\in$$\{u,i,r_{ui}\};$
$\text{With}\quad(Ecc_{x_{i}}-Ecc_{x_{i-1}})\geq$ 00 0
IV-B Usage of EAUC for model evaluation
@@
Moreover, since eccentricity bias leads to discriminatory outcomes when fairness is involved, EAUC also behaves as measure of fairness in dyadic regression tasks, being able to complement and synergize with classic fairness metrics. In this aspect, it is plausible that EAUC will become a reliable detector of an -until now- unknown source of unfairness in this type of task, i.e. eccentricity bias.
As an indicator of unfairness in dyadic regression, EAUC may be used to directly detect the presence of eccentricity-based individual or group unfairness [19], which can later be characterized using utility or parity-based fairness metrics [20]. Conversely, the inverse is also possible: while most group or individual unfairness are often attributed to representational or class-stereotypical biases in the data [21], they often cannot fully explain the source of fairness[22]; EAUC and its analysis may be used to better characterize the source of otherwise opaque unfairness phenomenons. One final property of EUAC is its ability to, unlike classic individual and group-based metrics modelling parity and utility [23], detect eccentricity-based unfairness that affects all entities (e.g. users) of the population as a whole, rather than only specific subgroups or individuals. As a result, EAUC can improve our ability to both detect and better characterize new and known unfairness in dyadic regression.